OpenStudy (anonymous):

evaluate the expression for x= 3 , 4-6 *(x-5) ??

OpenStudy (anonymous):

If you substitute 3 for x the equation will be 4-6*(3-5)

OpenStudy (anonymous):

4-6*(3-5)=16

OpenStudy (anonymous):

can someone show me ?

OpenStudy (anonymous):

how u got that answer ?

OpenStudy (anonymous):

do you know order of operations (PEMDAS)?

OpenStudy (anonymous):

when you have an expression like that it is often unclear what you need to do so you carry out PEMDAS... order of operations: P: parenthesis E: exponents MD: multiplication division AS: add subtract

OpenStudy (anonymous):

do P first... whatever is in the parenthesis..(3-5)=-2

OpenStudy (anonymous):

4-6*(3-5) 4-6*(-2) then you do multiplication next 4-(-12) then you do subtraction 4-(-12) = 16